# OpenFGL: A Comprehensive Benchmarks for Federated Graph Learning

### v1

@@ Abstract.
Federated graph learning (FGL) has emerged as a promising distributed training paradigm for graph neural networks across multiple local systems without direct data sharing. This approach is particularly beneficial in privacy-sensitive scenarios and offers a new perspective on addressing scalability challenges in large-scale graph learning. Despite the proliferation of FGL, the diverse motivations from practical applications, spanning various research backgrounds and experimental settings, pose a significant challenge to fair evaluation. To fill this gap, we propose OpenFGL, a unified benchmark designed for the primary FGL scenarios: Graph-FL and Subgraph-FL. Specifically, OpenFGL includes 38 graph datasets from 16 application domains, 8 federated data simulation strategies that emphasize graph properties, and 5 graph-based downstream tasks. Additionally, it offers 18 recently proposed SOTA FGL algorithms through a user-friendly API, enabling a thorough comparison and comprehensive evaluation of their effectiveness, robustness, and efficiency. Empirical results demonstrate the ability of FGL while also revealing its potential limitations, offering valuable insights for future exploration in this thriving field.
@@
Xunkai Li, Yinlin Zhu, Boyang Pang, Guochen Yan, Zening Li, Yeyu Yan, Zhengyu Wu, Wentao Zhang, Rong-Hua Li, Guoren Wang. PVLDB, 14(1): XXX-XXX, 2023.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Recently, graphs have emerged as effective tools for capturing intricate interactions among real-world entities, leading to their widespread applications. Based on this, we can translate various business applications from industrial scenarios into different graph-based downstream tasks from the machine learning perspective. To generate effective graph entity embeddings for downstream tasks, graph neural network (GNN) is a promising technology, which simultaneously encodes node features and structural insights, garnering significant attention. This paradigm has been widely validated, including node-level financial fraud detection (Hyun et al., 2023; Qiu, 2023), link-level recommendation (Li et al., 2022a; Yang et al., 2023b), and graph-level bioinformatics (Qu et al., 2023; Gao et al., 2023).
@@
Despite their effectiveness, privacy regulations, growing business competition, and scalability issues in billion-level graph learning have increased concerns around direct data sharing, posing challenges for centralized data storage and model training in a single machine. To address this issue, federated graph learning (FGL) is proposed, which aims to facilitate collaborative GNN training among multiple local systems. Existing FGL benchmarks, such as FS-G (Wang et al., 2022a) (Year: 2022) and FedGraphNN (He et al., 2021) (Year: 2021), offer valuable insights but still have the following limitations: (1) Datasets: Limited to few application domains (e.g., citation networks and recommendation systems). (2) Algorithms: Missing recent SOTA FGL methods (e.g., 8 methods in 2023, 10+ methods in 2024). (3) Experiments: Lack of federated data simulation strategies that account for graph properties, inadequate support for various graph-based downstream tasks, and limited evaluation perspectives. While the research prospects and enthusiasm for FGL are prominent and growing, the absence of a comprehensive benchmark for fair comparison impedes its development. Specifically, the diversity of graph-based downstream tasks (i.e., node, link, and graph), the unique graph characteristics (i.e., feature, label, and topology), and the complexity of FGL evaluation (i.e., effectiveness, robustness, and efficiency) collectively pose significant obstacles to achieving a comprehensive understanding of the current FGL landscape. Consequently, there is an emergency need to develop a standardized benchmark.
@@
In this paper, we propose OpenFGL, which integrates 2 commonly used FGL scenarios, 38 datasets in 16 application domains, 8 graph-specific distributed data simulation strategies, 18 recently proposed SOTA algorithms, and 5 graph-based downstream tasks. These components are implemented with a unified API to facilitate fair comparisons and future development in a user-friendly manner. Based on this foundation, we provide a comprehensive evaluation of existing FGL algorithms, drawing the following valuable insights. For Effectiveness we advocate for quantifying the statistics in distributed graphs to define the graph-based federated heterogeneity formally. For Robustness, to facilitate the practical deployment of existing FGL algorithms, we emphasize the significant potential of personalized, multi-client collaboration, and privacy-preserving techniques in addressing challenges such as data noise, low client participation, data sparsity, and generalization in complex applications. For Efficiency, considering the industry-scale datasets, we encourage FGL developers to prioritize algorithmic scalability and propose innovative federated collaborative paradigms that bring substantial benefits in improving efficiency. To further illustrate the advantages of our proposed OpenFGL compared to existing FGL benchmarks, we provide a clear description in Table 1.
@@
Our contributions. (1) Comprehensive Benchmark. We propose OpenFGL, which integrates the 2 most prevalent FGL scenarios with 38 publicly available datasets covering 5 downstream tasks. Based on this foundation, we thoroughly examine the unique properties of graphs and propose 8 practical distributed settings from the perspective of data heterogeneity. Meanwhile, we integrate 18 SOTA FGL algorithms and advocate 3 orthogonal evaluation perspectives to establish comprehensive baselines and systematic investigation of FGL algorithms. (2) Valuable Insights. Leveraging the user-friendly API integrated into OpenFGL, we conducted extensive empirical studies and derived 10 valuable conclusions (See Sec. 4.1-4.4). Building upon this, we provide 6 key insights from the perspectives of effectiveness, robustness, and efficiency, outlining promising research directions for future FGL studies (See Sec. 5). (3) Open-sourced Library and Detailed Repository. We develop an easy-to-use and open-source library to support ongoing FGL studies, allowing users to evaluate their FGL algorithms or datasets with ease. Additionally, we conduct a comprehensive review of existing FGL studies and release a beginner-friendly repository to facilitate the growth of the FGL community. The code and related tutorial are available at https://github.com/xkLi-Allen/OpenFGL.
@@ 2. Problem Statement
1. Receive Message. Each client initializes its local model with the one received from the server $\mathbf{W}^{T}_{k}\leftarrow\widehat{\mathbf{W}}^{T}$;
@@
Table 1. Comparison between OpenFGL and existing FGL benchmarks.
FGL Benchmarks Dataset Domain FGL Algorithms Federated Data Simulation Tasks Evaluations
FS-G (Wang et al., 2022a) 7 3 (GCFL, FedGNN, FedSage) Single-based Label, Topology 3 Simple Effectiveness and Efficiency
FedGraphNN (He et al., 2021) 6 0 (FL Implement of GNNs) LDA-based Feature, Label 3 Simple Effectiveness and Efficiency
OpenFGL (Ours) 16 18 Baselines (Year: 2021-2024)
Graph-orientated and Multiple-based
Feature, Label, Topology, Cross-domain
5 4 Comprehensive Modules (Table 2)
@@ 3. Benchmark Design
In this section, we offer detailed descriptions of OpenFGL, with an overview provided in Table 2. Specifically, we present a presentation from the perspectives of FGL application scenarios (Sec. 3.1), algorithms (Sec. 3.2), and experimental evaluations (Sec. 3.3).
@@ 3.1. Data-level FGL Scenarios
In OpenFGL, we focus on the following two most prevalent FGL scenarios: (1) Graph-FL. The growing integration with graph-based techniques and AI4Science applications, such as drug discovery, has motivated this scenario, in which clients consider graphs as the data samples and pursue collaborative training between clients to acquire powerful models while preserving data privacy. (2) Subgraph-FL. Realistic applications in this scenario include node-level fraud detection for financial security and link-level user-item interactions for recommendation, with data stored in a distributed manner. Clients treat their data as subgraphs of a larger and more comprehensive global graph, and focus on utilizing nodes and edges as data samples for training. Due to regulatory constraints and the data-sensitive nature, clients seek a collaborative training scheme to develop well-trained models without direct data sharing.
@@
Datasets. To comprehensively evaluate existing FGL algorithms in the above two real-world application-driven scenarios, we have compiled a substantial collection of public datasets from various domains. Specifically, Regarding Graph-FL scenario, we conduct experiments on the compounds networks (MUTAG, BZR, COX2, DHFR, PTC-MR, AIDS, NCI1, hERG) (Debnath et al., 1991; Sutherland et al., 2003; Helma et al., 2001; Riesen and Bunke, 2008; Wale et al., 2008; Gaulton et al., 2017), protein networks (ENZYMES, DD, PROTEINS) (Borgwardt et al., 2005; Dobson and Doig, 2003; Helma et al., 2001), collaboration network (COLLAB) (Leskovec et al., 2005), movie network (IMDB-B/M)  (Yanardag and Vishwanathan, 2015), super-pixel networks (MNISTSuperPixels) (Monti et al., 2017), and point cloud networks (ShapeNet) (Yi et al., 2016). As for Subgraph-FL scenario, we perform experiments on the citation networks (Cora, Citeseer, PubMed, FedDBLP, ogbn-arxiv) (Yang et al., 2016; Wang et al., 2022a; Hu et al., 2020b), co-purchase networks (Amazon-Computers, Amazon-Photo, ogbn-products) (Shchur et al., 2018; Hu et al., 2020b), and co-author networks (CS, Physics) (Shchur et al., 2018), wiki-page networks (Chameleon, Squirrel) (Pei et al., 2020; Platonov et al., 2023b), actor network (Actor) (Pei et al., 2020), game synthetic network (Minesweeper) (Platonov et al., 2023b), crowd-sourcing network (Tolokers) (Platonov et al., 2023b), article syntax network (Roman-empire) (Platonov et al., 2023b), rating network (Amazon-rating) (Platonov et al., 2023b), social network (Questions) (Platonov et al., 2023b), and point cloud networks (PCPNet, S3DIS) (Guerrero et al., 2018; Armeni et al., 2016).
@@
Table 3. The statistical information of the experimental Graph-FL datasets.
@@
ShapeNet 16,881 2616.2 KNN 3 50 40%/40%/40% Point Cloud Network
@@
Table 4. The statistical information of the experimental Subgraph-FL datasets.
@@
Actor 7,600 29,926 931 5 50/25%/25% Actor Network
Minesweeper 10,000 39,402 7 2 50/25%/25% Game Synthetic Network
Tolokers 11,758 519,000 10 2 50/25%/25% Crowd-sourcing Network
Roman-empire 22,662 32,927 300 18 50/25%/25% Article Syntax Network
Amazon-ratings 24,492 93,050 300 5 50/25%/25% Rating Network
Questions 48,921 153,540 301 2 50/25%/25% Social Network
@@
Simulation Strategies. In response to policy constraints on acquiring distributed graphs, we draw inspiration from federated learning in computer vision to simulate generalized federated scenarios by partitioning existing datasets into distributed subsets. This strategy is similar to recent FGL benchmarks (He et al., 2021; Wang et al., 2022a).
@@
In our proposed OpenFGL, we integrate 8 federated data simulation strategies driven by practical applications, in which we ensure the graph data distributed to each client exhibits similar patterns in feature, label, or topology while maintaining a controllable heterogeneity across clients. Specifically, for both Graph-FL and Subgraph-FL, we implement 3 simulation strategies widely used in graph-independent FL without graph-oriented designs, including Feature Distribution Skew, Label Distribution Skew, and Cross-Domain Data Skew. In the context of Graph-FL, we introduce a topology-oriented simulation strategy called Topology Shift, which distributes graphs based on degree distribution. The inspiration for this approach stems from key insights offered by recent FGL studies (Xie et al., 2021; Tan et al., 2023): in Graph-FL, the structure Non-iid resulting from topology shift is the primary challenge in collaborative optimization, where node features and labels appear less significant by comparison. As for Subgraph-FL, existing federated data simulation strategies predominantly utilize prevalent community detection algorithms such as Louvain (Blondel et al., 2008) and Metis (Karypis and Kumar, 1998) to identify clusters with dense intra-community connections. Then, these clusters’ nodes are subsequently allocated to local clients to construct corresponding induced subgraphs for partitioning. These strategies all operate under a common assumption: that the private data collected by each local agent in the real world contains dense internal connections but is loosely connected across clients. Despite their effectiveness, limitations persist in their application. This is because Subgraph-FL primarily focuses on node-level and link-level downstream tasks, where node profiles (i.e., node features and labels) are crucial. However, these strategies do not consider label distribution for federated data simulation. Hence, we introduce Metis-based Label Imbalance Split and Louvain-based Label Imbalance Split. These methods, refining the aforementioned strategies, carefully consider label distribution during cluster allocation to better simulate realistic and generalized distributed scenarios. We outline these strategies in Table 5, with descriptions provided below:
@@
Feature Distribution Skew is a graph-independent strategy to simulate feature distribution shifts. In OpenFGL, we utilize this approach to create more challenging and realistic scenarios for evaluating FGL algorithms. Specifically, we implement the following feature operations: (1) Adding Gaussian or Laplacian noise to introduce variability; (2) Applying scaling operations to simulate different magnitudes of features; (3) Employing mathematical transformations to further diversify the feature distributions.
@@
Label Distribution Skew is also a graph-independent simulation strategy. In our implementation, we use the $\alpha$-based Dirichlet distribution to create imbalanced label distributions across clients. This approach ensures varied and imbalanced label distributions, simulating real-world data scenarios. The $\alpha$ in Dirichlet distribution controls the concentration of probabilities across label classes, with larger values leading to more uniform distributions and smaller values resulting in sparser and more concentrated distributions.
@@
Cross Domain Data Skew is a fundamental challenge in FL, arising from the heterogeneous nature of data sources and collection methods across distributed databases. In OpenFGL, we simulate this scenario by evenly distributing multiple datasets among a predefined number of clients, maintaining a diverse representation.
@@
Topology Distribution Skew represents the strategy for partitioning graphs based on their topology properties in Graph-FL. This approach involves sorting the global graph according to specific characteristics, such as average node degree, and then distributing the resulting graphs to predefined clients. This ensures that the distribution of graph data among clients accurately reflects the underlying topological diversity of the original global dataset.
@@
Metis-based Label Imbalance Split is a new Subgraph-FL federated data simulation strategy introduced in this paper. The naive Metis-based Community Split lacks post-processing capabilities, leading to challenges in controlling subgraph heterogeneity among clients. In contrast, Metis+ enables predefined community partitioning, followed by clustering based on label distribution similarity, thereby consolidating similar communities under a single client.

### v2

# OpenFGL: A Comprehensive Benchmark for Federated Graph Learning

@@ Abstract.
Federated graph learning (FGL) is a promising distributed training paradigm for graph neural networks across multiple local systems without direct data sharing. This approach inherently involves large-scale distributed graph processing, which closely aligns with the challenges and research focuses of graph-based data systems. Despite the proliferation of FGL, the diverse motivations from real-world applications, spanning various research backgrounds and settings, pose a significant challenge to fair evaluation. To fill this gap, we propose OpenFGL, a unified benchmark designed for the primary FGL scenarios: Graph-FL and Subgraph-FL. Specifically, OpenFGL includes 42 graph datasets from 18 application domains, 8 federated data simulation strategies that emphasize different graph properties, and 5 graph-based downstream tasks. Additionally, it offers 18 recently proposed SOTA FGL algorithms through a user-friendly API, enabling a thorough comparison and comprehensive evaluation of their effectiveness, robustness, and efficiency. Our empirical results demonstrate the capabilities of FGL while also highlighting its potential limitations, providing valuable insights for future research in this growing field, particularly in fostering greater interdisciplinary collaboration between FGL and data systems.
@@
Xunkai Li, Yinlin Zhu, Boyang Pang, Guochen Yan, Zening Li, Yeyu Yan, Zhengyu Wu, Wentao Zhang, Rong-Hua Li, Guoren Wang. OpenFGL: A Comprehensive Benchmark for Federated Graph Learning. PVLDB, 14(1): XXX-XXX, 2023. doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Recently, graphs have emerged as effective tools for capturing intricate interactions among real-world entities, leading to their widespread applications. Based on this, we can translate various business applications from industrial scenarios into different graph-based downstream tasks from the machine learning perspective. To generate effective graph entity embeddings, graph neural networks (GNNs) utilize relational data stored from databases, encoding both node features and structural information for various data systems applications (Besta et al., 2022; Tian, 2023; Zhou et al., 2023). This paradigm has been widely validated, including node-level financial fraud detection (Hyun et al., 2023; Qiu, 2023), link-level recommendation (Li et al., 2022b; Yang et al., 2023b), and graph-level bioinformatics (Qu et al., 2023b; Gao et al., 2023).
@@
Despite their effectiveness, privacy regulations and scalability issues pose challenges to direct data sharing, complicating centralized model training (Yuan et al., 2021; Li et al., 2024a; Yuan et al., 2024). To address this challenge, federated graph learning (FGL) has been proposed to enable collaborative training across multiple local systems (Pan et al., 2023; Gu et al., 2023a; Zhang et al., 2024e; Yao et al., 2023; Wu et al., 2023), providing a novel distributed approach to graph-based data management (Sheth and Larson, 1990; Yang et al., 2019; Zhu et al., 2021; Antunes et al., 2022; Liu et al., 2022). Existing FGL benchmarks, such as FS-G (Wang et al., 2022a) (Year: 2022) and FedGraphNN (He et al., 2021) (Year: 2021), offer valuable insights but still have the following limitations: (1) Datasets: Limited to few application domains (e.g., citation networks and recommendation). (2) Algorithms: Missing recent SOTA FGL methods (e.g., 8 methods in 2023, 10+ methods in 2024). (3) Experiments: Lack of graph-oriented federated data simulation strategies, inadequate support for various graph-based downstream tasks, and limited evaluation perspectives. While the research prospects and enthusiasm for FGL are prominent and growing (Wu et al., 2024b; Liang et al., 2023; Wang et al., 2022c; Jiang et al., 2023), the absence of a comprehensive benchmark for fair comparison impedes its development. Specifically, the diversity of downstream tasks (i.e., node, link, and graph), the unique graph properties (i.e., feature, label, and topology), and the complexity of FGL evaluation (i.e., effectiveness, robustness, and efficiency) collectively pose significant obstacles to achieving a comprehensive understanding of the current FGL landscape. Consequently, there is an emergency need to develop a standardized benchmark.
@@
In this paper, we propose OpenFGL, which integrates 2 commonly used FGL scenarios, 42 datasets in 18 application domains, 8 graph-specific distributed data simulation strategies, 18 recently proposed SOTA algorithms, and 5 graph-based downstream tasks. These components are implemented with a unified API to facilitate fair comparisons and future development in a user-friendly manner. Based on this foundation, we provide a comprehensive evaluation of existing FGL algorithms, drawing the following valuable insights. For Effectiveness we advocate for quantifying the statistics in distributed graphs to define the graph-based federated heterogeneity formally. For Robustness, to facilitate the practical deployment of existing FGL algorithms, we emphasize the significant potential of personalized, multi-client collaboration, and privacy-preserving techniques in addressing challenges such as data noise, data sparsity, low client participation, and generalization in complex applications. For Efficiency, considering the industry-scale datasets, we encourage FGL developers to prioritize algorithmic scalability and propose innovative federated collaborative paradigms that bring substantial benefits in improving efficiency. To further illustrate the advantages of our proposed OpenFGL compared to existing FGL benchmarks, we provide a clear description in Table 1.
@@
Our contributions. (1) Comprehensive Benchmark. We propose OpenFGL, which integrates 2 scenarios with 42 publicly datasets. Based on this, we propose 8 practical distributed settings from the perspective of data heterogeneity. Meanwhile, we integrate 18 FGL algorithms and advocate 3 orthogonal evaluation perspectives to establish comprehensive baselines (See Table 1). We believe that FGL can inspire new research directions within the data systems community, particularly in scalable, privacy-preserving, and distributed data processing. (2) Valuable Insights. Leveraging the user-friendly API integrated into OpenFGL, we conducted extensive empirical studies and derived 10 valuable conclusions (See Sec. 4.1-Sec. 4.4). Building upon these findings, we provide 6 key insights from the perspectives of effectiveness, robustness, and efficiency, outlining promising research directions for the future FGL community (See Sec. 5). (3) Open-sourced Library and Detailed Repository. We develop an easy-to-use and open-source library to support ongoing FGL studies, allowing users to evaluate their algorithms or datasets with ease. Additionally, we conduct a comprehensive review of existing FGL studies and release a beginner-friendly repository to facilitate the growth of the FGL community. The code and related tutorial are available at https://github.com/xkLi-Allen/OpenFGL.
@@ 2. Problem Statement
1. Receive Message. Each client initializes its local model with the unified parameters from the server at the $t$-th round $\mathbf{W}^{T}_{k}\leftarrow\widehat{\mathbf{W}}^{T}$;
@@
Table 1. FGL benchmark comparison, where D.D. denotes dataset domain. Effe. and Effi. represent effectiveness and efficiency.
FGL Benchmarks D.D. FGL Algorithms Federated Data Simulation Tasks Evaluations Conclusions
FS-G (Wang et al., 2022a) 7 3 (Year: 2021) Label, Topology 3 Effe. and Effi. 5 (Effe.+Hyperparameter)
FedGraphNN (He et al., 2021) 6 0 (FL+GNNs) LDA-based Feature, Label 3 Effe. and Effi. 5 (Effe.+Effi.+Security)
OpenFGL (Ours) 18 18 (Year: 2021-2024)
Cross-domain and Graph-based
Feature, Label, Topology
5 Table 2
10 (Effe. + Robustness + Effi.)
and 6 Promising Directions
@@ 3.1. Data-level FGL Scenarios
In this section, we distinguish FGL scenarios based on the types of downstream tasks and the storage forms of local data at each client, categorizing them as Graph-FL, Subgraph-FL, and Node-FL. This serves as the guideline for proposing the concept of data-level FGL scenarios, emphasizing the data-centric description of real-world FGL applications. Notably, this section focuses on the pre-experiment preparation from a data perspective, whereas Sec. 3.3 emphasizes a more in-depth empirical analysis through a comprehensive evaluation across 3 orthogonal perspectives. In OpenFGL, we focus on the two prevalent FGL scenarios: (1) Graph-FL. The growing integration with graph-based techniques and AI4Science applications, such as drug discovery, has motivated this scenario, in which clients consider graphs as the data samples and pursue collaborative training between clients to acquire powerful models while preserving data privacy. (2) Subgraph-FL. Realistic applications in this scenario include node-level fraud detection for financial security and link-level user-item interactions for recommendation, with data stored in a distributed manner. Clients treat their data as subgraphs of a larger and more comprehensive global graph and focus on utilizing nodes and edges as data samples for training. Due to regulatory constraints, clients seek a collaborative training scheme to develop well-trained models without direct data sharing.
@@
Datasets. To comprehensively evaluate existing FGL algorithms, we have compiled a substantial collection of public datasets from various domains. Specifically, Regarding Graph-FL scenario, we conduct experiments on the compounds networks (MUTAG, BZR, COX2, DHFR, PTC-MR, AIDS, NCI1, hERG, ogbg-molhiv, ogbg-molpcba) (Debnath et al., 1991; Sutherland et al., 2003; Helma et al., 2001; Riesen and Bunke, 2008; Wale et al., 2008; Gaulton et al., 2017; Hu et al., 2020b), protein networks (ENZYMES, DD, PROTEINS, ogbg-ppa) (Borgwardt et al., 2005; Dobson and Doig, 2003; Helma et al., 2001; Hu et al., 2020b), collaboration network (COLLAB) (Leskovec et al., 2005), movie network (IMDB-B/M)  (Yanardag and Vishwanathan, 2015), super-pixel networks (MNISTSuperPixels) (Monti et al., 2017), point cloud (ShapeNet) (Yi et al., 2016), and syntax trees (ogbg-code2) (Hu et al., 2020b). As for Subgraph-FL scenario, we perform experiments on the citation networks (Cora, Citeseer, PubMed, FedDBLP, ogbn-arxiv) (Yang et al., 2016; Wang et al., 2022a; Hu et al., 2020b), co-purchase networks (Amazon-Computers, Amazon-Photo, ogbn-products) (Shchur et al., 2018; Hu et al., 2020b), and co-author networks (CS, Physics) (Shchur et al., 2018), wiki-page networks (Chameleon, Squirrel) (Pei et al., 2020; Platonov et al., 2023b), actor network (Actor) (Pei et al., 2020), game synthetic network (Minesweeper) (Platonov et al., 2023b), crowd-sourcing network (Tolokers) (Platonov et al., 2023b), syntax network (Roman-empire) (Platonov et al., 2023b), rating network (Amazon-rating) (Platonov et al., 2023b), social network (Questions) (Platonov et al., 2023b), and point cloud networks (PCPNet, S3DIS) (Guerrero et al., 2018; Armeni et al., 2016).
@@
Table 3. The statistical information of Graph-FL datasets.
@@
ogbg-molhiv 41,127 25.50 27.50 9 2 80%/10%/10% Compounds Network
ogbg-molpcba 437,929 26.00 28.10 9 2 80%/10%/10% Compounds Network
@@
ShapeNet 16,881 2616.20 KNN 3 50 40%/40%/40% Point Cloud Network
@@
ogbg-code2 452,741 125.20 124.20 4 275 90%/5%/5% Abstract Syntax Trees
Table 4. The statistical information of Subgraph-FL datasets.
@@
     Actor      7,600      29,926      931      5      50%/25%/25%      Actor Network
     Minesweeper      10,000      39,402      7      2      50%/25%/25%      Game Synthetic Network
     Tolokers      11,758      519,000      10      2      50%/25%/25%      Crowd-sourcing Network
     Roman-empire      22,662      32,927      300      18      50%/25%/25%      Article Syntax Network
     Amazon-ratings      24,492      93,050      300      5      50%/25%/25%      Rating Network
     Questions      48,921      153,540      301      2      50%/25%/25%      Social Network
@@
Simulation Strategies. In response to policy constraints on acquiring distributed graphs, we draw inspiration from federated learning in computer vision to simulate generalized federated scenarios by partitioning existing datasets into distributed subsets. This strategy is similar to recent FGL benchmarks (He et al., 2021; Wang et al., 2022a). In our proposed OpenFGL, we integrate 8 federated data simulation strategies driven by practical applications, in which we ensure the graph data distributed to each client exhibits similar patterns in feature, label, or topology while maintaining a controllable heterogeneity across clients. Specifically, for both Graph-FL and Subgraph-FL, we implement 3 simulation strategies widely used in graph-independent FL. In the context of Graph-FL, we introduce a topology-oriented simulation strategy called Topology Shift, which distributes graphs based on degree distribution. The inspiration for this approach stems from key insights offered by recent FGL studies (Xie et al., 2021; Tan et al., 2023): in Graph-FL, the structure Non-iid resulting from topology shift is the primary challenge in collaborative optimization, where node features and labels appear less significant by comparison. As for Subgraph-FL, existing strategies predominantly utilize community detection algorithms such as Louvain (Blondel et al., 2008) and Metis (Karypis and Kumar, 1998) to identify clusters with dense intra-community connections. Then, these clusters’ nodes are subsequently allocated to local clients to construct corresponding induced subgraphs for partitioning. These strategies all operate under a common assumption: that the private data collected by each local agent in the real world contains dense internal connections but is loosely connected across clients (Hsiao, 1992; Kamel and Kamel, 1992; Li et al., 2020a, 2021b). Despite their effectiveness, limitations persist in their application. This is because Subgraph-FL primarily focuses on node-level and link-level tasks, where node profiles (i.e., node features and labels) are crucial. However, these strategies do not consider label distribution for federated data simulation. Hence, we introduce Metis-based Label Imbalance Split and Louvain-based Label Imbalance Split. These methods, refining the aforementioned strategies, carefully consider label distribution during cluster allocation to better simulate realistic and generalized distributed scenarios. We outline these strategies in Table 5, with descriptions provided below:
@@
Feature Distribution Skew is a graph-independent strategy to simulate feature distribution shifts (Wang et al., 2022c; Jiang et al., 2023). In OpenFGL, we utilize this approach to create more challenging and realistic scenarios for evaluating FGL algorithms (Li et al., 2022a). Specifically, we implement various feature operations: (1) Adding Gaussian or Laplacian noise to introduce variability; (2) Applying scaling operations to simulate different magnitudes of features; (3) Employing mathematical transformations to further diversify the feature distributions.
@@
Label Distribution Skew is a graph-independent strategy (Wu et al., 2024b; Yao et al., 2023). In our implementation, we use the $\alpha$-based Dirichlet distribution to create imbalanced label distributions across clients (Blei et al., 2003). This approach ensures varied and imbalanced label distributions, simulating real-world data scenarios. The $\alpha$ in Dirichlet distribution controls the concentration of probabilities across label classes, with larger values leading to more uniform distributions and smaller values resulting in sparser and more concentrated distributions.
@@
Cross Domain Data Skew is a fundamental challenge in FL, arising from the heterogeneous nature of data sources and collection methods across distributed databases (Yang et al., 2019). In OpenFGL, we simulate this scenario by evenly distributing multiple datasets among a predefined number of clients, maintaining a diverse representation.
@@
Topology Distribution Skew represents the strategy for partitioning graphs based on their topology properties in Graph-FL (Xie et al., 2021; Tan et al., 2023). This approach involves sorting the global graph according to specific characteristics, such as average node degree, and then distributing the resulting graphs to predefined clients. This ensures that the distribution of graph data among clients accurately reflects the underlying topological diversity of the original global dataset.
@@
Metis-based Label Imbalance Split is a new Subgraph-FL data simulation strategy introduced in this paper. The naive Metis-based Community Split lacks post-processing capabilities, leading to challenges in controlling subgraph heterogeneity among clients. In contrast, our approach enables predefined community partitioning, followed by clustering based on label distribution similarity, thereby consolidating similar communities under a single client.